OpenStudy (bananas1234):

Find the constant of variation for the relationship f(x)= 50x. x f(x) 5 50

Mehek (mehek14):

what will stay the same?

Mehek (mehek14):

@Bananas1234 ?

Mehek (mehek14):

f(x) depends on the value 5 and 50 are just a numbers the value of x will never change for the equation

OpenStudy (bananas1234):

i see

OpenStudy (bananas1234):

I think that makes sence

Mehek (mehek14):

so what do you think the answer is?

OpenStudy (bananas1234):

A?

Mehek (mehek14):

yes

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

constant of variation is not x

OpenStudy (bananas1234):

ok

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=k~\cdot x }\) is the equation for the direct variation. \(\large\color{black}{ \displaystyle x }\) is the independent variable (it is whatever you plug in for x). \(\large\color{black}{ \displaystyle y }\) is the dependent variable (it depends on what value you are going to plug in for x). \(\large\color{black}{ \displaystyle k }\) is the constant that relates y and x. In this case, whatever you plug it for x, the y will be k times greater than that.

OpenStudy (bananas1234):

So do i have to slove for K?

OpenStudy (solomonzelman):

You don't need to "solve for k" just compare the (abstract) form y=kx to your equations

OpenStudy (solomonzelman):

equation*

OpenStudy (bananas1234):

50?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=k~\cdot x }\) formula. \(\large\color{black}{ \displaystyle y=50~\cdot x }\) your's

OpenStudy (solomonzelman):

if you have any questions, please ask.

OpenStudy (bananas1234):

So if K = 50 would that mean the Constant would be 50?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

Also, "constant" (by definition) is "a number" (not a variable) but any real number.

OpenStudy (bananas1234):

Thank you. So for future refrence i just compare my given equation to Y = K times X?

OpenStudy (solomonzelman):

well, for this question, or questions that involve different numbers but have the same format, then yes.

OpenStudy (bananas1234):

Thanks

OpenStudy (solomonzelman):

or, you can say this....

OpenStudy (solomonzelman):

``` "constant of variation" is a number that relates the 2 variables. What are the variables here? Well, they are x and y. What does the 50 do? Any number that you plug in for x (call it a), will be multiplied times 50 to find y (and that is regardless of what the value of a is). So k is relating y and x, by saying that any y-value is greater than x-value [as you plug in any number for x]. ```

OpenStudy (solomonzelman):

hope this made sense just now.

OpenStudy (bananas1234):

I think so, math is kinda hard for me. I will save that info. thanks.

OpenStudy (solomonzelman):

Alright... not that it is easy for anyone (well, on person's own level). You are welcome!